Subject: APO 3.5-m Users Committee minutes 8/9/04

From: Bruce Gillespie

Submitted: Wed, 11 Aug 2004 13:37:26 -0600

Message number: 824 (previous: 823, next: 825 up: Index)

					APO 3.5-m Users Committee Phonecon, 8/9/04

Attending:  Ed Turner, Jon Holtzman, Michael Strauss, Russet McMillan, 
Karl Glazebrook, Rene Walterbos, Don York, Bruce Balick, Bruce 
Gillespie, Suzanne Hawley, John Barentine

Absent:  Al Harper, Jim Green, John Bally

Minutes taken by Bruce Gillespie

**********************************


Engineering shutdown status (including the extension):

Bruce G reported that nearly all of the shutdown tasks have been 
completed or are in a state of good progress.  All the instruments have 
had their annual service except for GRIM2 and the Echelle, which has 
been pumped but the detector will not be cooled until next week.  GRIM2 
servicing will be done next week.  The damaged mirror covers have been 
repaired, and the telescope reassembly has begun.  This week, we are 
working on the air cooling hose for the new NA2 guider, which required 
using a bigger diameter plenum than earlier planned, so some mechanical 
modifications are being made.  All in all, it looks like we may be 
ready for science as predicted on 26 August unless an unexpected 
show-stopper pops up.  Jon Holtzman asked if we were doing additional 
things during the shutdown since there is more time than originally 
scheduled, such as TUI work.  The answer is no, for the most part.  
Russell has been working on TUI, but Craig has been heavily involved in 
the software for the new NA2 guider.  Also, with a more limited labor 
pool than in previous years, we are barely able to accomplish the 
routine annual housekeeping jobs during this shutdown, especially since 
some of the staff is shared with SDSS which has been in its own 
engineering shutdown for the past month.  Unless more manpower can be 
found and applied to our future shutdowns, making any significant 
improvements to the telescope will continue to be problematic.  Don 
York asked about the new NA2 guider, and in the following discussion is 
was agreed that we should consider sparing the new camera and making 
sure that our other guider cameras had ready spares, e.g., for the 
Echelle guider.

**********************************


NICFPS status and plans:

Jon Barentine has been seconded to the NIC-FPS instrument team for the 
past few weeks to help with final assembly and lab tests.  He reported 
that the instrument is getting close to being finished, and he expects 
an initial delivery to APO and late September.

According to John, the optics have been integrated and cold-cycled, as 
have all the mechanical components.  They have had "First Lab Light" 
and the pinhole image on the detector is 3 pixels fwhm, or smaller--the 
shape of the image is OK.  The mechanical cold tests all worked fine, 
but they haven't yet tested the IR etalon itself because they are 
waiting for parts.  All the optical filters are on hand, and will be 
cold tested shortly.  After the initial fit tests on the telescope in 
September, it is planned to take the instrument back to Boulder for 
some rework, if necessary.  Then the instrument would return to APO in 
October for full commissioning tests, which are expected to take up to 
10 nights.  The ARC user community will be soon asked to propose 
science commissioning observations; the NIC-FPS team will make explicit 
what level of involvement in data collection and reduction will be 
provided for these tests.  Ed Turner mentioned that having John 
Barentine working with the NIC-FPS instrument team has proven to be 
invaluable, and is a lesson we should remember for future projects.  He 
also said that the allocation for telescope time during Q4 assumes that 
10 nights will be devoted to NIC-FPS commissioning.

**********************************


UVa/triplespec situation:

Suzanne Hawley started the discussion by recapping the history and 
current status of our involvement and discussions with the University 
of Virginia.  John Wilson and Mike Skrutskie brought their CorMASS 
instrument to APO in April, and had a very successful observing run in 
early May on targets mostly proposed by ARC scientists.  This led to 
informal discussions between several UVa astronomers and Ed, Suzanne, 
and Bruce G about the possibilities of keeping CorMASS at APO, ARC's 
possible interest in having UVa build us a medium-resolution IR 
spectrometer (called Triplespec, with R~3000, which UVa has plans to 
build up to three instruments for other telescopes), and whether there 
was a mutual interest in exploring how UVa and ARC could collaborate in 
general.

Ed, Suzanne, and Bruce G plan a visit to UVa in late August to further 
discuss these possibilities with their faculty and staff.  Ed mentioned 
several potential benefits of a collaborative arrangement: we get a 
very desirable instrument, we develop a liaison with a strong telescope 
and instrumentation group, and potentially could get much-needed 
engineering support for the telescope.  What we need soon is for the 
Users Committee members to poll their users to see what level of 
support they have for Triplespec and possible affiliation between ARC 
and UVa. The committee was largely supportive of the idea and 
encouraged pursuing further discussions with UVa.  It was also 
mentioned that the ARC BoG is also discussing/considering possible 
arrangements with UVa.  Don York pointed out that the data reduction 
software for Triplespec data would likely be very difficult to write 
and use, and that we should make sure this is part of any arrangement 
to have UVa build Triplespec for us.

**********************************


Sloan SN follow-up project in 4Q2004

Ed and Don said that a proposal would be submitted to the UChicago TAC 
for Q4 to test the proposed SDSS Supernovae extension project.  During 
dark and grey time, the 2.5-m will find supernovae and the 3.5-m will 
make followup observations of their light curves at a cadence of every 
2 to 3 nights.  The 3.5-m would be scheduled for these observations in 
advance, on the assumption that the supernovae will be readily found.  
If approved, this test program will create a new scheduling constraint 
for Ed, making it difficult if not impossible to schedule other dark or 
grey runs than need several consecutive nights on the telescope.  On 
the other hand, if the tests are successful it will demonstrate the 
synergistic power of the 2.5-m/3.5-m working in tandem on a very 
important area of astrophysical research, lending support to the 
current SDSS extension projects proposal which is under review by the 
NSF.

**********************************


Last month's minutes were approved without comment.

Next phonecon will be on Monday, Sept. 13, 2004, at 11:30 AM Eastern 
Time

 


APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO
APO  This is message 824 in the apo35-general archive. You can find
APO  the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO  To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO  To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO